Moe is 6 times Curley's age. If the difference inthei ages is 20 years, how old is Moe?
Found 2 solutions by ewatrrr, iyah98:
Answer by ewatrrr(24785)

Hi,
Let x be Curley's age. Then Moe's age is 6x
.
Question States
6x - x = 20
.
5x = 20
x = 4yr (Curley's age)
.
Moe is 6*4yr or 24yrs old.

Answer by iyah98(2)

Let M be the age of Moe and C be the age of Curley
Equation base on the given statement
M = 6C
M = C + 20
Simplified
C + 20 = 6C
5C = 20
C = 4
THEREFORE Moe's age is
M = 6 x 4
M = 24